# Script2Screen: Supporting Dialogue Scriptwriting with Interactive Audiovisual Generation

### v1

@@ Abstract.
Figure 1. The Script2Screen interface, captured from a writing session. Script2Screen allows screenwriters to seamlessly visualize and hear their scenes. Writers can select a workspace from the Scenes navigation bar (A), where the corresponding script loads into the Script Editor (B). After the script is processed, the Screen Preview (C) offers a live 3D animation with synced audio, complete with customizable camera angles and background settings (D). The Timeline View (E) provides a streamlined, interactive breakdown of each dialogue segment, showcasing emotions, speaker cues, and playback controls in dialogue cards (F). Script2Screen delivers a fully immersive WYSIWYG experience, making it an ideal tool for crafting scripts, brainstorming ideas, or generating dynamic storyboards ready for downstream applications.
[…]
Pipeline of Script2Screen.
@@ 3.1. Gap between Scriptwriting and Audiovisual Production
On the other hand, the visual artists’ perspective reveals another dimension of this disconnect. P4, a technical animator responsible for visualizing scripts through animation, expressed the challenges of not receiving visual input from writers. P4 highlighted that when writers do not provide clear visual references, it becomes the responsibility of animators or directors to interpret and fill in these gaps:
”It would be nice if, in pre-production, they could provide reference videos of what they want things to look like, because otherwise, I have to be the one to look for them.” This reliance on production team members to make creative decisions adds a layer of complexity. While writers may feel they lose control over their work, animators and directors may feel burdened with interpreting and visualizing the script in a way that aligns with the writer’s vision. This disconnect often results in discrepancies between the writer’s original intent and the final product.
@@ 7.1. Study Procedure
The study was conducted remotely or in person, depending on participants’ preferences. Each session lasted 75 minutes and was designed to compare our system, Script2Screen, with a baseline common writing tool, Google Docs, presented in a randomized order. Initially, we asked users to choose any editor for scriptwriting. We found that the first few expert participants all chose Google Docs over professional scriptwriting editors. They explained that professional tools such as Final Draft [^3] https://www.finaldraft.com/ could be excessive for the tasks we designed, while Google Docs was sufficient and also allowed for live presentation of results. Observing this trend, we decided to use Google Docs as the baseline to ensure consistency across participants (both experts and novices).
@@ 7.3. Study Results
Figure 8. Comparison of average ratings between baseline and Script2Screen across the Creative Support Index. Error bars represent the standard deviation. Statistical significance between the two groups is indicated with asterisks: * ($p<0.05$), **∗ ∗ ($p<0.01$), and ***∗ ∗ ∗ ($p<0.001$). Brackets highlight significant differences, showing strong evidence of improvements in several metrics for the Script2Screen tool compared to the Baseline in almost all categories.

### v2

@@ Abstract.
Figure 1. The Script2Screen interface, captured from a writing session. Script2Screen allows screenwriters to seamlessly visualize and hear their scenes. Writers can select a workspace from the Scenes navigation bar (A), where the corresponding script loads into the Script Editor (B). After the script is processed, the Screen Preview (C) offers a live 3D animation with synced audio, complete with customizable camera angles and background settings (D). The Timeline View (E) provides a streamlined, interactive breakdown of each dialogue segment, showcasing emotions, speaker cues, and playback controls in dialogue cards (F). Script2Screen delivers a fully immersive WYSIWYG experience, making it an ideal tool for crafting scripts, brainstorming ideas, or generating dynamic storyboards ready for downstream applications.:
@@ 3.1. Gap between Scriptwriting and Audiovisual Production
On the other hand, the visual artists’ perspective reveals another dimension of this disconnect. P4, a technical animator responsible for visualizing scripts through animation, expressed the challenges of not receiving visual input from writers.
P4 highlighted that when writers do not provide clear visual references, it becomes the responsibility of animators or directors to interpret and fill in these gaps: ”It would be nice if, in pre-production, they could provide reference videos of what they want things to look like, because otherwise, I have to be the one to look for them.” This reliance on production team members to make creative decisions adds a layer of complexity. While writers may feel they lose control over their work, animators and directors may feel burdened with interpreting and visualizing the script in a way that aligns with the writer’s vision. This disconnect often results in discrepancies between the writer’s original intent and the final product.
@@ 6.4. System Architecture
6.5. Pipeline Design Choice vs. End-to-End Video Generative Models
At the time of this study, state-of-the-art audio-visual generative models such as Sora2 and Veo3 had not yet been released. Existing video generation methods also had several limitations, which motivated our pipeline-based design.
Generation Length.
Most end-to-end video models only support a few seconds of coherent output. In contrast, our pipeline can co-generate audio and animation of arbitrary length, limited only by the input script and compute resources. This makes it more suitable for extended dialogue scenes and interactive storytelling.
Iterative Control.
End-to-end generation couples visual quality with narrative control, which makes small edits difficult. For example, changing a line of text or adjusting a gesture often requires regenerating the entire video. Our modular pipeline supports iteration at different stages (script, audio, or animation), enabling a more flexible and user-driven workflow.
Modularity and Extensibility.
Each component of the pipeline (LLM parsing, TTS, gesture generation, and animation retargeting) is implemented as a separate module. This makes it possible to improve or replace individual parts, and to integrate new models as they become available. The modular design ensures that the system can continue to evolve beyond the current limits of end-to-end video generation.
@@ 7.1. Study Procedure
The study was conducted remotely or in person, depending on participants’ preferences. Each session lasted 75 minutes and was designed to compare our system, Script2Screen, with a baseline common writing tool, Google Docs, presented in a randomized order. Initially, we asked users to choose any editor for scriptwriting. We found that the first few expert participants all chose Google Docs over professional scriptwriting editors for short writing tasks in our study. They explained that professional tools such as Final Draft [^3] https://www.finaldraft.com/ could be excessive for the tasks we designed, while Google Docs was sufficient and also allowed for live presentation of results. Observing this trend, we decided to use Google Docs as the baseline to ensure consistency across participants (both experts and novices).
@@ 7.3. Study Results
Figure 8. Comparison of average ratings between baseline and Script2Screen across the Creative Support Index. Error bars represent the standard deviation. Statistical significance between the two groups is indicated with asterisks: * ($p<0.05$), ** ($p<0.01$), and *** ($p<0.001$). Brackets highlight significant differences, showing strong evidence of improvements in several metrics for the Script2Screen tool compared to the Baseline in almost all categories.
@@ 10. Conclusion
11. GenAI Usage Disclosure
In accordance with the ACM Policy on the Use of Generative AI in Publication, we disclose that generative AI (GenAI) tools were used both as part of the research and during the preparation of this manuscript.
  • Research Tooling: Our system integrates GenAI models directly. An audio generative model (2024) for audio generation, a gesture generative model (2023) for animation, and large language model (2023) served as the backbone for dispatching and analyzing natural language.
  • Authoring Support: We used ChatGPT for limited assistance in (i) writing plotting code for Figures 8 and 9, (ii) identifying and summarizing related works, and (iii) improving transitions and flow in the User Study section.
All substantive research contributions, study design, implementation, data analysis, and final interpretations were carried out by the authors.